A laser beam is transmitted with a "width" of 0.00200 degree. What is the diameter d of a spot of the beam on an object 258,000km distant?
Answer by stanbon(75887)

A laser beam is transmitted with a "width" of 0.00200 degree. What is the diameter d of a spot of the beam on an object 258,000km distant?
-------------
Draw the picture (an isosceles triangle)
Draw a perpendicular bisector forming two right triangles at the diameter line.
---------
Pick one of the triangles; it has angle 0.001 degree and base 258,000 km
Find the altitude (a)
tan(0.001) = a/258,000
a = 258000*tan(0.001)
a = 4.50 km
That is the radius of the beam.
-----
Diameter of the beam = 2*4.5 = 9 km
============
Cheers,
Stan H.